# AB-UPT: Scaling Neural CFD Surrogates for High-Fidelity Automotive Aerodynamics Simulations via Anchored-Branched Universal Physics Transformers

### v2

@@ Abstract
Recent advances in neural surrogate modeling offer the potential for transformative innovations in applications such as automotive aerodynamics. Yet, industrial-scale problems often involve volumetric meshes with cell counts reaching the 100 millions, presenting major scalability challenges. Complex geometries further complicate modeling through intricate surface-volume interactions, while quantities such as vorticity are highly nonlinear and must satisfy strict divergence-free constraints. To address these requirements, we introduce  Anchored-Branched Universal Physics Transformers (AB-UPT) as a novel modeling scheme for building neural surrogates for  computational fluid dynamics (CFD) simulations. AB-UPT is designed to: (i) decouple geometry encoding and prediction tasks via multi-branch operators; (ii) enable scalability to high-resolution outputs via neural simulation in a low-dimensional latent space, coupled with anchored neural field decoders to predict high-fidelity outputs; (iii) enforce physics consistency by a novel divergence-free formulation. We show that AB-UPT yields state-of-the-art predictive accuracy of surface and volume fields on automotive CFD simulations ranging from 33 thousand up to 150 million mesh cells. Furthermore, our anchored neural field architecture enables the enforcement of hard physical constraints on the physics predictions without degradation in performance, exemplified by modeling divergence-free vorticity fields. Notably, the proposed models can be trained on a single GPU in less than a day and predict industry-standard surface and volume fields within seconds. Additionally, we show that the flexible design of our method enables neural simulation from a computer-aided design geometry alone, omitting the need for costly CFD meshing procedures.
@@ 1 Introduction
Figure 1: AB-UPT is a neural surrogate model trained to jointly model surface and volume variables of automotive CFD simulations with $>100$M simulation mesh cells. It obtains state-of-the-art surface and volume predictions (left), accurately models drag and lift coefficients (center) all on a single GPU (right).
@@
In recent years, deep neural network-based surrogates have emerged as a computationally efficient alternative in science and engineering (Thuerey et al., 2021; Zhang et al., 2023; Brunton et al., 2020), impacting, e. g., weather forecasting (Pathak et al., 2022; Bi et al., 2023; Lam et al., 2023; Nguyen et al., 2023; Bodnar et al., 2025), protein folding (Jumper et al., 2021; Abramson et al., 2024), or material design (Merchant et al., 2023; Zeni et al., 2025; Yang et al., 2024). In many of these scientific disciplines, the Transformer architecture (Vaswani et al., 2017) has become a central backbone of the recent breakthroughs. Inherently linked to Transformers is the notion of scale, both in terms of trainable parameters of the model and the amount of data needed to train those scaled models. However, the notion of scale is ill-defined in the field of neural surrogate modeling for science and engineering, mainly due to the lack of availability of high-volume training data. Hence, we start this work by posing the following question:
@@ 2 Preliminaries and related work
Table 1: Conceptual overview of different CFD solvers. Numeric integration schemes, while accurate, are slow and require specialized meshes to converge, with several orders of magnitude variations in fidelity and runtime. Neural surrogates struggle to be trained on the full-resolution mesh, which makes them slow to train or requires multi-GPU setups for million-scale meshes. Next to that, those methods require computation units that favor speed over accuracy. We introduce AB-UPT, a new method incorporating anchor tokens. AB-UPT unlocks linear problem scaling, which enables training of neural surrogates on a single GPU in less than a day, even on million-scale meshes. This efficiency also allows us to leverage expressive neural computation units such as the Transformer. Furthermore, our multi-branch design enables simulation with any input mesh, uses reduced latent modeling of physical dynamics. Finally, anchor attention acts as a neural field, enabling an efficient divergence-free vorticity field formulation. *We hypothesize that neural surrogates can produce results with an accuracy similar to RANS when trained on data from HRLES. However, such comparisons are difficult to make as there are no public datasets available which have paired RANS/HRLES results for the same large-scale geometries.
@@ 2.2 Transformer blocks for building neural surrogates
Many neural surrogate models are formulated in the neural operator learning paradigm (Lu et al., 2021a; Li et al., 2020; 2021; Seidman et al., 2022; Alkin et al., 2024a; b). In this framework, neural networks represent operators that map between Banach spaces $\mathcal{I}$ and $\mathcal{O}$ of functions defined on compact domains $\mathcal{X}$ and $\mathcal{Y}$, i. e., ${\mathcal{G}}:\mathcal{I}\rightarrow\mathcal{O}$. Neural operators provide continuous outputs that remain consistent across varying input sampling resolutions and patterns. Training a neural operator involves constructing a dataset of input-output function pairs evaluated at discrete spatial locations.
@@
Cross-attention is a variant of scaled dot-product attention where a second latent representation $\mathbf{Z}_{KV}\in\mathbb{R}^{M\times d}$ is used as input to $\mathbf{K}(\cdot)$ and $\mathbf{V}(\cdot)$. Intuitively, this formulation incorporates information from $\mathbf{Z}_{KV}$ into $\mathbf{Z}$. It is commonly used to enable interactions between different modeling components (e. g., Vaswani et al. (2017)), data types (e. g., Alayrac et al. (2022)), or to compress/expand representations (e. g., Jaegle et al. (2021)). Roughly speaking, the first two use-cases use $N\approx M$ whereas the third use-case uses $N\ll M$ (compression) or $M\ll N$ (expansion).
@@ 2.3 Related work
In this section, we will discuss work related to neural surrogate modeling for aerodynamics.
@@
Training neural surrogates directly on a complete simulation mesh can be computationally expensive and, in some cases, even infeasible. To address this issue, compression of the input mesh into a reduced latent space can be employed. Next, a decoder (often functioning as a point-wise decoder) needs to be able to reconstruct the output mesh conditioned on the compressed latent representation and single query points. To satisfy this requirement, the decoder must function as a conditional neural field. Such approaches are introduced in AROMA (Serrano et al., 2024), CViT (Wang et al., 2024b), Knigge et al. (2024), GOAT Wen et al. (2025), LNO (Wang & Wang, 2024), and UPT (Alkin et al., 2024a). Moreover, UPT additionally introduces a patch-embedding analogue for compressing general geometries via supernode pooling.
@@
Methods like Reg-DGCNN (Wang et al., 2019; Elrefaie et al., 2024a; b), Graph-UNet Gao & Ji (2019), PointNet (Qi et al., 2017a), and transformer models with linear complexity, such as FactFormer (Li et al., 2024), GNOT (Hao et al., 2023), ONO (Xiao et al., 2023), OFormer (Li et al., 2023a) or Transolver (Wu et al., 2024) have constant compute complexity w.r.t. the number of input points. This allows such models to process the unreduced simulation mesh at high computation costs, as shown in Transolver++ (Luo et al., 2025), which demonstrates processing meshes with 2.5 million cells via sequence parallelism on four GPUs. Although technically feasible, such approaches also become infeasible at larger scales. For example, linearly extrapolating the reported GPU requirements of Transolver++ would necessitate roughly 256 GPUs to process a simulation mesh with 140 million cells. In contrast, AB-UPT can process simulations of 140 million cells on a single GPU.
@@
Decoupling encoding and decoder components is often beneficial for tasks like predicting 3D flow fields directly from geometric inputs (see e. g., GINO (Li et al., 2023b), GOAT Wen et al. (2025), LNO Wang et al. (2024a), OFormer (Li et al., 2023a), and UPT). Additionally, a decoupled latent space representation allows for scalable decoding to a large number of output mesh cells since such models can cache encoded latent states and decode output queries in parallel. Moreover, decoupling the encoder and decoder also allows for independent input and output meshes. This means the model does not require the same simulation mesh used for ground truth output as its encoder input during training (i.e., the encoder can be simulation mesh independent), providing greater flexibility in data handling.
@@
Physics-Informed Neural Networks (PINN) (Raissi et al., 2019) are a popular approach for learning neural networks that enforce the governing equations by adding the residuals of the partial differential equation (PDE) and boundary conditions as soft penalty terms in the loss function. It has been shown that these methods require second-order optimization methods to learn from the otherwise conflicting signal in the data and the soft constraints Wang et al. (2025). In this work, we instead use the conditional field property of our model to directly incorporate physical properties into the architecture itself as a hard constraint. Related approaches (i.e., (Lu et al., 2021b; Richter-Powell et al., 2022)) are much less common and have so far not been scaled to the massive model, data, and problem scales considered in this work.
@@ 3.1 Multi-branch architecture
We discretize the input of the surface and volume branches by turning the simulation mesh into sets of 3D point clouds. Specifically, we select $N^{{s}}$ points from the surface and $N^{{v}}$ points in the 3D volume, i. e., $\mathbf{X}^{{s}}\in\mathbb{R}^{N^{{s}}\times 3}$ and $\mathbf{X}^{{v}}\in\mathbb{R}^{N^{{v}}\times 3}$. To embed the 3D coordinates, we use the Transformer positional encoding as in the geometry branch. Then, we encode these embedded points by using two multi-layer perceptron (MLP) encoders, $\mathcal{E}^{{s}}$ and $\mathcal{E}^{{v}}$. We employ separate MLP encoders for surface coordinates ($\mathcal{E}^{{s}}$) compared to volume coordinates ($\mathcal{E}^{{v}}$) to provide explicit contextual information, designating which points reside on the surface and within the volume. To summarize, the surface and volume encoder are defined as follows:
@@
The final decoder block consist of $L$ self-attention blocks that do not share parameters and hence operate independently. The query, key, and value representations all originate from the previous layer, and, therefore, there is no interaction anymore between the two branches. This design allows for branch-specific modeling, meaning each branch can refine the output predictions independently, without interaction with the other branches. Finally, a single linear projection is applied as a decoder to generate the final output predictions for each branch.
@@ 3.2 Anchor attention
To mitigate this problem, we introduce the concept of anchor attention, which splits the discretized input mesh (i. e., the point cloud) into two parts: (i) a small subset $\mathbf{A}$ of $M$ anchor tokens is uniformly sampled from the $N$ input tokens and used to derive the key matrix $\mathbf{K}(\mathbf{A})$ and value matrix $\mathbf{V}(\mathbf{A})$ for the attention mechanism; (ii) the query matrix $\mathbf{Q}(\mathbf{Z})$, which is derived from all $N$ input tokens, then cross-attend to the anchor key and value matrices. Intuitively, this formulation splits the $N$ input tokens into $M$ anchor tokens (used as queries, keys and values in the attention) and $N-M$ query tokens (used only as queries in the attention). It is equivalent to self-attention within the $M$ anchor tokens and cross-attention between $N-M$ query tokens and the $M$ anchor tokens. Formally, anchor attention is defined as follows:
@@
Importantly, the $N-M$ query tokens only need to compute attention with a much smaller set of $M$ key-value pairs (derived from the anchor tokens) compared to full self-attention, which computes all $N^{2}$ pairs. Anchor attention therefore reduces the computational complexity of the attention layers from $\mathcal{O}\left(N^{2}\right)$ to $\mathcal{O}\left(M^{2}+(N-M)M\right)=\mathcal{O}\left(MN\right)$, where $M^{2}$ is the self-attention between the $M$ anchor tokens and $(N-M)M$ is the cross-attention from $N-M$ query tokens to $M$ anchor tokens. This is crucial because surrogate models often need to be evaluated at a large number of input points (e. g., to obtain accurate drag/lift coefficient predictions as shown in Figure 1). Additionally, since there are no interactions between query tokens, arbitrary many query tokens can be used per forward pass. Therefore, query tokens can be processed in chunks, enabling constant memory consumption by using a constant amount of queries. When processing all $N$ input tokens at once (end-to-end inference), memory complexity of anchor attention is $\mathcal{O}(N)$ (we use an efficient attention implementation (Dao et al., 2022)). However, as $N$ can be $>$ 100M, it would still run into out-of-memory errors on a single GPU. By chunking the $N-M$ query tokens into chunks of size $C$, memory complexity reduces to $\mathcal{O}(M+C)$, i.e., constant memory consumption. We refer to this methodology as chunked inference (c. f. inference memory in Figure 1). Moreover, in chunked inference, the key-value pairs of anchor tokens can be cached when calculating them in the first chunked forward pass to omit recalculating them for subsequent chunks (anchor tokens remain constant for all chunks). However, as $M\ll N$, this is often negligible.
Finally, by treating the anchors as conditioning set for the queries, anchor attention turns the architecture into a conditional field model: the field values at any query location are predicted from the query location itself and the context provided by the anchors, but independently from any other query location. This has two important benefits. First, this enables creating predictions at arbitrary query positions, which allows AB-UPT to predict at arbitrary resolution. Many simulation insights can be obtained from much lower resolutions than necessary for numeric simulation (e.g., drag/lift coefficients in Figure 6). Second, it allows us to impose physics-based constraints directly into the model predictions, such as a divergence-free condition for vorticity (see Section 3.3).

### v3

@@ Abstract
Recent advances in neural surrogate modeling offer the potential for transformative innovations in applications such as automotive aerodynamics. Yet, industrial-scale problems often involve volumetric meshes with cell counts reaching 100 million, presenting major scalability challenges. Complex geometries further complicate modeling through intricate surface-volume interactions, while quantities such as vorticity are highly nonlinear and must satisfy strict divergence-free constraints. To address these requirements, we introduce  Anchored-Branched Universal Physics Transformers (AB-UPT) as a novel modeling scheme for building neural surrogates for  computational fluid dynamics (CFD) simulations. AB-UPT is designed to: (i) decouple geometry encoding and prediction tasks via multi-branch operators; (ii) enable scalability to high-resolution outputs via neural simulation in a low-dimensional latent space, coupled with anchored neural field decoders to predict high-fidelity outputs; (iii) enforce physics consistency by a novel divergence-free formulation. We show that AB-UPT yields state-of-the-art predictive accuracy of surface and volume fields on automotive CFD simulations ranging from 33 thousand up to 150 million mesh cells. Furthermore, our anchored neural field architecture enables the enforcement of hard physical constraints on the physics predictions without degradation in performance, exemplified by modeling divergence-free vorticity fields. Notably, the proposed models can be trained on a single GPU in less than a day and predict industry-standard surface and volume fields within seconds. Additionally, we show that the flexible design of our method enables neural simulation from a computer-aided design geometry alone, omitting the need for costly CFD meshing procedures.
@@ 1 Introduction
Figure 1: AB-UPT is a neural surrogate model trained to jointly model surface and volume variables of automotive CFD simulations with $>100$M simulation mesh cells. It obtains state-of-the-art surface and volume predictions (left), accurately models drag and lift coefficients (center), all on a single GPU (right).
@@
In recent years, deep neural network-based surrogates have emerged as a computationally efficient alternative in science and engineering (Brunton et al., 2020; Thuerey et al., 2021; Zhang et al., 2023), impacting, e. g., weather forecasting (Pathak et al., 2022; Bi et al., 2023; Lam et al., 2023; Nguyen et al., 2023; Bodnar et al., 2025), protein folding (Jumper et al., 2021; Abramson et al., 2024), or material design (Merchant et al., 2023; Zeni et al., 2025; Yang et al., 2024). In many of these scientific disciplines, the Transformer architecture (Vaswani et al., 2017) has become a central backbone of the recent breakthroughs. Inherently linked to Transformers is the notion of scale, both in terms of trainable parameters of the model and the amount of data needed to train those scaled models. However, the notion of scale is ill-defined in the field of neural surrogate modeling for science and engineering, mainly due to the lack of availability of high-volume training data. Hence, we start this work by posing the following question:
@@ 2 Preliminaries and related work
Table 1: Conceptual overview of different CFD solvers. Numeric integration schemes, while accurate, are slow and require specialized meshes to converge, with several orders of magnitude variations in fidelity and runtime. Neural surrogates struggle to be trained on the full-resolution mesh, which makes them slow to train or requires multi-GPU setups for million-scale meshes. Next to that, those methods require computation units that favor speed over accuracy. We introduce AB-UPT, a new method incorporating anchor tokens. AB-UPT unlocks linear problem scaling, which enables training of neural surrogates on a single GPU in less than a day, even on million-scale meshes. This efficiency also allows us to leverage expressive neural computation units such as the Transformer. Furthermore, our multi-branch design enables simulation with any input mesh, uses reduced latent modeling of physical dynamics. Finally, anchor attention acts as a neural field, enabling an efficient divergence-free vorticity field formulation. *It’s difficult to make comparisons between RANS simulations and models that are trained on HRLES simulations, since RANS simulations cannot resolve certain phenomena which HRLES can resolve. One example is strong separation.
@@ 2.2 Transformer blocks for building neural surrogates
Many neural surrogate models are formulated in the neural operator learning paradigm (Li et al., 2020; 2021; Lu et al., 2021; Seidman et al., 2022; Alkin et al., 2024a; b). In this framework, neural networks represent operators that map between Banach spaces $\mathcal{I}$ and $\mathcal{O}$ of functions defined on compact domains $\mathcal{X}$ and $\mathcal{Y}$, i. e., ${\mathcal{G}}:\mathcal{I}\rightarrow\mathcal{O}$. Neural operators provide continuous outputs that remain consistent across varying input sampling resolutions and patterns. Training a neural operator involves constructing a dataset of input-output function pairs evaluated at discrete spatial locations.
@@
Cross-attention is a variant of scaled dot-product attention where a second latent representation $\mathbf{Z}_{KV}\in\mathbb{R}^{M\times d}$ is used as input to $\mathbf{K}(\cdot)$ and $\mathbf{V}(\cdot)$. Intuitively, this formulation incorporates information from $\mathbf{Z}_{KV}$ into $\mathbf{Z}$. It is commonly used to enable interactions between different modeling components (e. g., (Vaswani et al., 2017)), data types (e. g., (Alayrac et al., 2022)), or to compress/expand representations (e. g., (Jaegle et al., 2021)). Roughly speaking, the first two use-cases use $N\approx M$ whereas the third use-case uses $N\ll M$ (compression) or $M\ll N$ (expansion).
@@ 2.3 Related work
In this section, we will discuss work related to neural surrogate modeling and hard physical constraints in neural networks.
@@
Training neural surrogates directly on a complete simulation mesh can be computationally expensive and, in some cases, even infeasible. To address this issue, compression of the input mesh into a reduced latent space can be employed. Next, a decoder (often functioning as a point-wise decoder) needs to be able to reconstruct the output mesh conditioned on the compressed latent representation and single query points. To satisfy this requirement, the decoder must function as a conditional neural field. Such approaches are introduced in AROMA (Serrano et al., 2024), CViT (Wang et al., 2024b), Knigge et al. (2024), GOAT (Wen et al., 2025), LNO (Wang & Wang, 2024), and UPT (Alkin et al., 2024a). Moreover, UPT additionally introduces a patch-embedding analogue for compressing general geometries via supernode pooling.
@@
Methods like Reg-DGCNN (Wang et al., 2019; Elrefaie et al., 2024a; b), Graph-UNet (Gao & Ji, 2019), PointNet (Qi et al., 2017a), and Transformer models with linear complexity, such as Erwin (Zhdanov et al., 2025), FactFormer (Li et al., 2024a), GNOT (Hao et al., 2023), ONO (Xiao et al., 2023), OFormer (Li et al., 2023a) or Transolver (Wu et al., 2024) have constant compute complexity w.r.t. the number of input points. This allows such models to process the unreduced simulation mesh at high computation costs, as shown in Transolver++ (Luo et al., 2025), which demonstrates processing meshes with 2.5 million cells via sequence parallelism on four GPUs. Although technically feasible, such approaches also become infeasible at larger scales. For example, linearly extrapolating the reported GPU requirements of Transolver++ would necessitate roughly 256 GPUs to process a simulation mesh with 140 million cells. In contrast, AB-UPT can process simulations of 140 million cells on a single GPU.
@@
Decoupling encoding and decoder components is often beneficial for tasks like predicting 3D flow fields directly from geometric inputs (see e. g., GINO (Li et al., 2023b), GOAT (Wen et al., 2025), LNO (Wang et al., 2024a), OFormer (Li et al., 2023a), and UPT). Additionally, a decoupled latent space representation allows for scalable decoding to a large number of output mesh cells since such models can cache encoded latent states and decode output queries in parallel. Moreover, decoupling the encoder and decoder also allows for independent input and output meshes. This means the model does not require the same simulation mesh used for ground truth output as its encoder input during training (i.e., the encoder can be simulation mesh independent), providing greater flexibility in data handling.
@@
A popular approach to embed physical principles into neural networks is to use the governing equations as a soft constraint, penalizing the partial differential equation (PDE) residuals in the loss function (Raissi et al., 2019; Li et al., 2024b). This encourages but does not strictly enforce physical consistency, and often requires second-order methods to alleviate optimization challenges (Wang et al., 2025). An alternative is to enforce hard constraints, for instance by projecting the output of an unconstrained model onto a physically feasible solution space (Négiar et al., 2023; Hansen et al., 2023; Chalapathi et al., 2024; Utkarsh et al., 2025). In contrast, our work enforces physical principles such as the divergence-free nature of the vorticity field as a hard constraint by construction. Our approach is closely related to the work by Richter-Powell et al. (2022) and Liu et al. (2024), who construct divergence-free networks using differential operators. However, to our knowledge, none of the aforementioned physics-informed approaches have been scaled to the massive model, data, and problem sizes considered in this work.
@@ 3.1 Multi-branch architecture
We discretize the input of the surface and volume branches by turning the simulation mesh into sets of 3D point clouds. Specifically, we select $N^{{s}}$ points from the surface and $N^{{v}}$ points in the 3D volume, i. e., $\mathbf{X}^{{s}}\in\mathbb{R}^{N^{{s}}\times 3}$ and $\mathbf{X}^{{v}}\in\mathbb{R}^{N^{{v}}\times 3}$. To embed the 3D coordinates, we use the Transformer positional encoding as in the geometry branch. Then, we encode these embedded points by using two multi-layer perceptron (MLP) encoders, $\mathcal{E}^{{s}}$ and $\mathcal{E}^{{v}}$. We employ separate MLP encoders for surface coordinates ($\mathcal{E}^{{s}}$) compared to volume coordinates ($\mathcal{E}^{{v}}$) to provide explicit contextual information, designating which points reside on the surface and within the volume. To summarize, the surface and volume encoders are defined as follows:
@@
The final decoder block consists of $L$ self-attention blocks that do not share parameters and hence operate independently. The query, key, and value representations all originate from the previous layer, and, therefore, there is no interaction anymore between the two branches. This design allows for branch-specific modeling, meaning each branch can refine the output predictions independently, without interaction with the other branches. Finally, a single linear projection is applied as a decoder to generate the final output predictions for each branch.
@@ 3.2 Anchor attention
To mitigate this problem, we introduce the concept of anchor attention, which splits the discretized input mesh (i. e., the point cloud) into two parts: (i) a small subset $\mathbf{A}$ of $M$ anchor tokens is uniformly sampled from the $N$ input tokens and used to derive the key matrix $\mathbf{K}(\mathbf{A})$ and value matrix $\mathbf{V}(\mathbf{A})$ for the attention mechanism; (ii) the query matrix $\mathbf{Q}(\mathbf{Z})$, which is derived from all $N$ input tokens, then cross-attend to the anchor key and value matrices. Intuitively, this formulation splits the $N$ input tokens into $M$ anchor tokens (used as queries, keys, and values in the attention) and $N-M$ query tokens (used only as queries in the attention). It is equivalent to self-attention within the $M$ anchor tokens and cross-attention between $N-M$ query tokens and the $M$ anchor tokens. Formally, anchor attention is defined as follows:
@@
Importantly, the $N-M$ query tokens only need to compute attention with a much smaller set of $M$ key-value pairs (derived from the anchor tokens) compared to full self-attention, which computes all $N^{2}$ pairs. Anchor attention therefore reduces the computational complexity of the attention layers from $\mathcal{O}\left(N^{2}\right)$ to $\mathcal{O}\left(M^{2}+(N-M)M\right)=\mathcal{O}\left(MN\right)$, where $M^{2}$ is the self-attention between the $M$ anchor tokens and $(N-M)M$ is the cross-attention from $N-M$ query tokens to $M$ anchor tokens. This is crucial because surrogate models often need to be evaluated at a large number of input points (e. g., to obtain accurate drag/lift coefficient predictions as shown in Figure 1).
Additionally, since there are no interactions between query tokens, an arbitrary number of query tokens can be used per forward pass. Therefore, query tokens can be processed in chunks, enabling constant memory consumption by using a constant number of queries. When processing all $N$ input tokens at once (end-to-end inference), the memory complexity of anchor attention is $\mathcal{O}(N)$ (we use an efficient attention implementation (Dao et al., 2022)). However, as $N$ can be $>$ 100 million, it would still run into out-of-memory errors on a single GPU. By chunking the $N-M$ query tokens into chunks of size $C$, memory complexity reduces to $\mathcal{O}(M+C)$, i.e., constant memory consumption. We refer to this methodology as chunked inference (c. f. inference memory in Figure 1). Moreover, in chunked inference, the key-value pairs of anchor tokens can be cached when calculating them in the first chunked forward pass to omit recalculating them for subsequent chunks (anchor tokens remain constant for all chunks). However, as $M\ll N$, this is often negligible.
Finally, by treating the anchors as conditioning set for the queries, anchor attention turns the architecture into a conditional field model: the field values at any query location are predicted from the query location itself and the context provided by the anchors, but independently from any other query location. This has two important benefits. First, this enables creating predictions at arbitrary query positions, which allows AB-UPT to predict at arbitrary resolution. Many simulation insights can be obtained from much lower resolutions than necessary for numeric simulation (e.g., drag/lift coefficients in Figure 6). Second, it allows us to impose physics-based constraints directly into the model predictions, such as a divergence-free condition for vorticity (see Section 3.3). In the context of Neural Processes, the concept of conditioning the output prediction on a (small) subset of context has been explored in (Feng et al., 2023; Ashman et al., 2024).